### v2

@@ 1 Introduction
Financial markets are characterized by their complex, non-linear behaviors, posing significant challenges for modeling and prediction. While traditional approaches to market analysis rely heavily on statistical tools and Euclidean assumptions, these methods often fail to account for deeper geometric structures inherent in the data. A novel perspective can be gained by incorporating concepts from differential geometry, particularly the classification of 2-manifolds into three geometries: spherical, Euclidean, and hyperbolic.Thurston1997
@@
The geometrization theorem, which underpins this classification, provides a foundational framework for understanding how spaces can be decomposed into simpler geometric components,Perelman2002. This insight, central to modern differential geometry, suggests that many complex systems, including financial markets, may exhibit characteristics aligning with one or more of these geometric types. As in physics and network science, the widespread application of geometric methods and their potential in financial market analysis is actively explored under active research,CapponiLehalle2023; SimonianFabozzi2019; LopezDePrado2018AFML; LopezDePrado2020MLAM; NoguerAlonso2021FinEAS. Recent studies explore the application of manifold learning techniques to financial market analysis and time series forecasting. These methods aim to extract low-dimensional representations of complex, high-dimensional data, revealing intrinsic structures and patterns (Y2017Nonlinear; G2018Nonlinear; Y2020Manifold). In Jansen2023 several trading practice applications and concepts are outlined, as well. Manifold learning approaches have been used for early warning systems in financial markets (Y2017Nonlinear; G2018Nonlinear), phase space reconstruction of financial time series (Y2014Manifold), and visualization of market states (Y2020Manifold). Researchers have also developed novel algorithms, such as information metric-based manifold learning (Y2017Nonlinear) and kernel entropy manifold learning (Y2014kernel), to improve the accuracy of financial analysis and prediction. Beyond finance, manifold learning techniques have been applied to various time series analysis tasks, including electroencephalography signal analysis (P2018Multivariate) and forecasting of high-dimensional time series (P2021Time), demonstrating their versatility in capturing complex dynamics across different domains.Perelman2002
@@
Brownian motion, a stochastic process traditionally studied in Euclidean spaces, is a powerful tool for modeling random dynamics. Extending Brownian motion to spherical and hyperbolic geometries (Hsu2002; IkedaWatanabe1989) enables the representation of processes influenced by curvature and global topological features. Studies have explored diffusion on spheres (Gomez2021geometrical; M2000Brownian) and hyperbolic spaces (L2007Hyperbolic, Hsu2002StochasticAO), as well as fractional Brownian motion in both geometries (J2005Spherical). The effects of curvature on Brownian motion have been investigated using Riemann normal coordinates (Castro-Villarreal, 2010) and the Smoluchowski equation (Pavel2014Intrinsic). Research has shown that positive curvature slows diffusion, while negative curvature accelerates it (Pavel2010Brownian). The concept of Brownian motion has been generalized to metric spaces of constant curvature (S1981Brownian). Additionally, studies have examined the hydrodynamics of curved membranes and their impact on particulate mobility (Henle2010Hydrodynamics), providing insights into diffusion on biological structures like vesicles and membrane tethers.Hsu2002; IkedaWatanabe1989
@@
Financial markets exhibit complex, nonlinear dynamics that may be better understood through non-Euclidean geometric frameworks (E2006Geometry; Ilinski1999Gauge;Ilinski2001; Emami2021; KellerRessel2021). These markets can be modeled using projective geometry, fiber bundles, and fractal geometry to capture their intricate structures and behaviors (E2006Geometry; Ilinski1999Gauge; B2004MIS; Lipton2001FX; Lipton2018SelectedWorks). Researchers have developed methods to incorporate non-Euclidean geometric information into filtering and machine learning algorithms, improving their accuracy in financial applications (Anastasis2017Geometric). Network geometry measures, such as discrete Ricci curvatures, can be used to analyze market instability and systemic risk (Samal2021Network). The concept of ”dark volatility,” a hidden factor influencing market behavior, has been explored using Einstein warped product manifolds (Pincak2023possible). These geometric approaches provide powerful tools for modeling asset behavior, forecasting volatility, and evaluating investment risks with greater precision (B2004MIS; Humphrey2011Financial).(Ilinski2001; KellerRessel2021; Emami2021)
@@ 3.1 Ambient formulation and curvature drift
Brownian motion constrained to $M$ is the Stratonovich SDE (Hsu2002)
@@
where $B_{t}$ is standard Brownian motion in $\mathbb{R}^{3}$. In Ito form this becomes (Hsu2002)
@@ 3.2 Explicit SDEs by geometry
    The metric is diagonal: $g_{\theta\theta}=(R+r\cos\varphi)^{2}$, $g_{\varphi\varphi}=r^{2}$,doCarmoRiemannian, hence $g^{\theta\theta}=(R+r\cos\varphi)^{-2}$, $g^{\varphi\varphi}=r^{-2}$ and $\sqrt{|g|}=r(R+r\cos\varphi)$. The Ito SDE realizing generator $\tfrac{1}{2}\Delta_{g}$ is
@@
Then $g_{uu}=E(u)=a^{2}\sinh^{2}u+c^{2}\cosh^{2}u$, $g_{vv}=G(u)=a^{2}\cosh^{2}u$,doCarmoRiemannian, so $g^{uu}=1/E$, $g^{vv}=1/G$, $\sqrt{|g|}=\sqrt{E(u)G(u)}=a\cosh u\,\sqrt{E(u)}$. The Ito SDE (generator $\tfrac{1}{2}\Delta_{g}$) reads
@@ 3.3 Nonlinear forecasting via machine learning regressors
A Random Forest (breiman2001random) constructs an ensemble of $B$ regression trees $\{f_{b}(\cdot)\}_{b=1}^{B}$, each trained on a bootstrap resample of $\mathcal{D}$ and a random subset of predictors. The forecast corresponds to the ensemble mean:
@@ 3.4 Logarithmic and exponential mappings
Let $T_{x}M$ denote the tangent plane at $x\in M$.Pennec2018; doCarmoRiemannian
@@
See Pennec2018; doCarmoRiemannian for closed forms.
Torus $T^{2}(R,r)$ (angle chart).doCarmoCurves
@@
Hyperbolic (hyperboloid chart).doCarmoRiemannian
@@
Karcher mean and tangent PCA.Karcher1977; Pennec2018
@@ 3.5 Data-driven estimation of manifold parameters
A key ingredient of our pipeline is that the geometry is not treated as fixed; instead, its intrinsic parameters are inferred on the fly from the observed 3D path $\{x_{t}=(x_{t},y_{t},z_{t})\}_{t\leq t^{\star}}$. This subsection formalizes the estimation steps we implement for the sphere, torus and one–sheeted hyperboloid, matching the code used in our experiments.CazalsPouget2005; CohenSteiner2006
@@
For spherical segments, we first compute a Karcher (Fréchet) mean $\widehat{\mu}\in S^{2}$ of the points $\{x_{t}/\|x_{t}\|\}$ by iterating $\mu\leftarrow\exp_{\mu}\!\big(\tfrac{1}{n}\sum_{t}\log_{\mu}(x_{t})\big)$ until convergence;Karcher1977
@@
In practice, for stability we use $\widehat{R}$ from (22) while the instantaneous tube radius is proxied by the latest $\rho_{t^{\star}}$, which helps track slow deformations of the tube thickness in non–stationary segments. doCarmoCurves
@@
On this chart, the log map again uses shortest arc wrapping on $v$ and simple differencing on $u$; the exp map re–adds those increments around the base point. doCarmoRiemannian
@@ 3.6 Local Gaussian Curvature Based Geometry Inference
At each time $t$ we consider a window $\mathcal{W}_{t}=\{X_{s}:s\in[t_{0},t]\}$ ( expanding) with $|\mathcal{W}_{t}|\geq m_{0}$ points and compute a local quadratic Monge patch fit,CazalsPouget2005; CohenSteiner2006,
@@
where $W_{t}=\mathrm{diag}(\,w_{1},\dots,w_{|\mathcal{W}_{t}|}\,)$ is optional exponential weighting (recent samples weighted more), with $w_{i}\propto e^{-\alpha(t-s_{i})}$. Given $\hat{a}_{t},\hat{b}_{t},\hat{c}_{t}$, the local Gaussian curvature at time $t$ is (Monge gauge approximation),CazalsPouget2005,
@@
Towards validating that the method approximates satisfactorily the underlying Gaussian curvature, Fig. 2 recovers the expected signatures on benchmarks—-$K>0$ on $S^{2}$, mixed $K$ on $T^{2}$, and $K<0$ on $H^{2}$—and yields plausible, intermittent curvature on the finance path; “uniformly sampled” means draws from each surface’s Riemannian (area) measure (density $\propto dA$).doCarmoRiemannian
@@
We form a Takens embedding, Takens1981, over the window (Takens1981) $\mathcal{W}_{t}$:
@@
with delay $\tau$ and embedding dimension $m$. We compute Vietoris Rips persistent homology of $\mathcal{T}_{t}$ and count persistent 1-cycles (Betti $1$ features),EdelsbrunnerHarer2010; Ghrist2008. A torus satisfies $\beta_{1}=2$ (and $\beta_{2}=1$),EdelsbrunnerHarer2010, so we flag
@@
with persistence threshold $\epsilon$ calibrated to the scale of the point cloud,Gidea2018; Ismail2022; Arvanitis2024. In the current context, we implement the basic $\beta_{1}$–based detector over sliding windows.

### v3

@@ 1 Introduction
Financial markets are characterized by their complex, non-linear behaviors, posing significant challenges for modeling and prediction. While traditional approaches to market analysis rely heavily on statistical tools and Euclidean assumptions, these methods often fail to account for deeper geometric structures inherent in the data. A novel perspective can be gained by incorporating concepts from differential geometry, particularly the classification of 2-manifolds into three geometries: spherical, Euclidean, and hyperbolic.Thurston (1997)
@@
The geometrization theorem, which underpins this classification, provides a foundational framework for understanding how spaces can be decomposed into simpler geometric components,Perelman (2002). This insight, central to modern differential geometry, suggests that many complex systems, including financial markets, may exhibit characteristics aligning with one or more of these geometric types. As in physics and network science, the widespread application of geometric methods and their potential in financial market analysis is actively explored under active research,Capponi and Lehalle (2023); Simonian and Fabozzi (2019); López de Prado (2018, 2020); Gutiérrez-Fandiño et al. (2021). Recent studies explore the application of manifold learning techniques to financial market analysis and time series forecasting. These methods aim to extract low-dimensional representations of complex, high-dimensional data, revealing intrinsic structures and patterns (Huang et al. (2017); Tzagkarakis (2018); Huang (2020)). In Jansen (2023) several trading practice applications and concepts are outlined, as well. Manifold learning approaches have been used for early warning systems in financial markets (Huang et al. (2017); Tzagkarakis (2018)), phase space reconstruction of financial time series (Huang and Kou (2014b)), and visualization of market states (Huang (2020)). Researchers have also developed novel algorithms, such as information metric-based manifold learning (Huang et al. (2017)) and kernel entropy manifold learning (Huang and Kou (2014a)), to improve the accuracy of financial analysis and prediction. Beyond finance, manifold learning techniques have been applied to various time series analysis tasks, including electroencephalography signal analysis (Rodrigues et al. (2018)) and forecasting of high-dimensional time series (Papaioannou et al. (2022)), demonstrating their versatility in capturing complex dynamics across different domains.Perelman (2002)
@@
Brownian motion, a stochastic process traditionally studied in Euclidean spaces, is a powerful tool for modeling random dynamics. Extending Brownian motion to spherical and hyperbolic geometries (Hsu (2002a); Ikeda and Watanabe (1989)) enables the representation of processes influenced by curvature and global topological features. Studies have explored diffusion on spheres (Gómez and Sevilla (2021); Krishna et al. (2000)) and hyperbolic spaces (Lao and Orsingher (2007), Hsu (2002b)), as well as fractional Brownian motion in both geometries (Istas (2005)). The effects of curvature on Brownian motion have been investigated using Riemann normal coordinates (Castro-Villarreal, 2010) and the Smoluchowski equation (Castro-Villarreal (2014)). Research has shown that positive curvature slows diffusion, while negative curvature accelerates it (Castro-Villarreal (2010)). The concept of Brownian motion has been generalized to metric spaces of constant curvature (Takenaka et al. (1981)). Additionally, studies have examined the hydrodynamics of curved membranes and their impact on particulate mobility (Henle and Levine (2010)), providing insights into diffusion on biological structures like vesicles and membrane tethers.Hsu (2002a); Ikeda and Watanabe (1989)
@@
Financial markets exhibit complex, nonlinear dynamics that may be better understood through non-Euclidean geometric frameworks (Piotrowski and Sładkowski (2006); Ilinski (1999);Ilinski (2001); Emami (2021); Keller-Ressel and Nargang (2021)). These markets can be modeled using projective geometry, fiber bundles, and fractal geometry to capture their intricate structures and behaviors (Piotrowski and Sładkowski (2006); Ilinski (1999); Mandelbrot and Hudson (2004); Lipton (2001, 2018)). Researchers have developed methods to incorporate non-Euclidean geometric information into filtering and machine learning algorithms, improving their accuracy in financial applications (Kratsios and Hyndman (2017)). Network geometry measures, such as discrete Ricci curvatures, can be used to analyze market instability and systemic risk (Samal et al. (2021)). The concept of ”dark volatility,” a hidden factor influencing market behavior, has been explored using Einstein warped product manifolds (Pinčák et al. (2023)). These geometric approaches provide powerful tools for modeling asset behavior, forecasting volatility, and evaluating investment risks with greater precision (Mandelbrot and Hudson (2004); Humphrey and Wong (2011)).(Ilinski (2001); Keller-Ressel and Nargang (2021); Emami (2021))
@@
2 Concept and Scope: From Linear Orthogonality to Geometric Structure
This study bridges the gap between linear dimensionality reduction and non-linear geometric modeling. Our approach rests on three conceptual pillars: the critical distinction between orthogonality and independence in PCA, the operational necessity of identifying a low-dimensional manifold via the Uniformization Theorem, and the rigorous construction of tradable eigenportfolios.
2.1 Dimensionality Reduction and Tradable Eigenportfolios
We employ an expanding window Principal Component Analysis (PCA) to extract the latent drivers of the market structure. Let $\mathbf{R}_{t}\in\mathbb{R}^{N}$ denote the vector of standardized log-returns for $N$ assets at time $t$. The empirical covariance matrix $\Sigma_{t}$ is decomposed via Spectral Decomposition:
$\Sigma_{t}=\mathbf{V}_{t}\mathbf{\Lambda}_{t}\mathbf{V}_{t}^{\top}$ (1)
where $\mathbf{V}_{t}$ is the matrix of eigenvectors (loadings) and $\mathbf{\Lambda}_{t}$ the diagonal matrix of eigenvalues.
To ensure the resulting portfolios are tradable and free from look-ahead bias, we construct the weights for the eigenportfolios at time $t$ as derived from the eigenvectors computed at time t1t-1. Thus, the return $F_{k,t}$ of the $k$-th eigenportfolio is realized as:
$F_{k,t}=\mathbf{R}_{t}\cdot\mathbf{v}_{k,t-1}$ (2)
This lag ensures that the portfolio composition is known at the close of t1t-1, making the strategy executable in a realistic trading environment Avellaneda and Lee (2010); Jolliffe (2002).
We retain the first three principal components ($k=1,2,3$), constructing a 3-dimensional embedding $\mathbf{X}_{t}=(F_{1,t},F_{2,t},F_{3,t})^{\top}\in\mathbb{R}^{3}$. Descriptive analysis of our dataset (2005–2025) confirms that this 3-dimensional projection captures, on average, 67.5% of the total market variance, 5.
It is crucial to note that while PCA imposes orthogonality (linear uncorrelatedness, $\mathbb{E}[F_{i}F_{j}]=0$ for $i\neq j$), it does not imply statistical independence ($\mathbb{P}(F_{i},F_{j})\neq\mathbb{P}(F_{i})\mathbb{P}(F_{j})$), unless the underlying data is jointly Gaussian. Financial time series exhibit ”stylized facts” such as volatility clustering and tail dependence Cont (2001), which imply significant non-linear coupling between uncorrelated modes. We hypothesize that this residual non-linear dependency manifests as the intrinsic curvature of the manifold, distinguishing the market’s true geometry from a flat Euclidean space.
2.2 The Uniformization Conjecture: A Parity of Dimensions
Our restriction to a 3-dimensional embedding is a deliberate topological choice rooted in the classification of surfaces. By modeling the market state as a trajectory on a 2-manifold embedded in $\mathbb{R}^{3}$, we utilize the Uniformization Theorem Abikoff (1981); Poincaré (1907). This fundamental result asserts that every simply connected Riemann surface is conformally equivalent to one of three canonical geometries:
  • Spherical ($S^{2}$): Positive curvature ($K>0$), implying compact, recurrent dynamics.
  • Euclidean ($\mathbb{R}^{2}$): Zero curvature ($K=0$), implying flat, random-walk dynamics.
  • Hyperbolic ($H^{2}$): Negative curvature ($K<0$), implying divergent or saddle-path dynamics.
We further include the Torus ($T^{2}$), which, although locally Euclidean, possesses a non-trivial topology ($S^{1}\times S^{1}$) essential for capturing cyclical market phases.
Extending the model to include just one additional principal component (4 PCs) would require modeling a 3-manifold. Under the Thurston-Perelman Geometrization Theorem Thurston (1997); Perelman (2002), the classification of 3-manifolds explodes in complexity, requiring decomposition into eight distinct geometries (e.g., $S^{3},\mathbb{H}^{3},\text{Nil},\text{Sol}$). By limiting our scope to 3 PCs, we maintain a tractable geometric framework where the curvature sign provides clear, actionable market signals. However, we recognize that capturing the full spectrum of market variance may eventually require this higher-dimensional leap; consequently, extending our framework to the full eight Thurston geometries remains the primary objective of our future research.
2.3 Curvature as an Alpha Signal
The superior performance of the Geometry-Informed Model (GIM) over ”flat” Euclidean baselines can be attributed to the correct specification of the transport dynamics. In a Euclidean framework (e.g., standard VAR), the expected path is a straight line in the tangent space, assuming zero acceleration.
However, if the market manifold is curved, a linear forecast systematically drifts off the manifold surface. The GIM corrects this by introducing a geometry-induced drift term derived from the mean curvature vector $\mathbf{H}$:
$d\mathbf{X}_{t}=\mathbf{P}(\mathbf{X}_{t})\circ d\mathbf{B}_{t}\approx\mathbf{P}(\mathbf{X}_{t})d\mathbf{B}_{t}-\frac{1}{2}\mathbf{H}(\mathbf{X}_{t})dt$ (3)
where $\mathbf{P}(\mathbf{X}_{t})$ projects the noise onto the tangent space Hsu (2002a). The term $\frac{1}{2}\mathbf{H}$ acts as a ”restoring force,” pulling the trajectory along the manifold’s geodesic. For example, in a spherical regime, this term enforces mean reversion naturally, whereas a flat model would predict a continued trend. By explicitly modeling this curvature, GIM anticipates non-linear turning points that linear models misinterpret as noise.
2.4 Macroeconomic Intuition: The Latent Triad
Beyond statistical variance and geometric tractability, the decision to target a 3-dimensional embedding finds a robust justification in macroeconomic theory. Financial asset prices are, in the long run, discounted claims on the real economy. Consequently, the dimensionality of the market’s factor space should mirror the dimensionality of the underlying macroeconomic state vector.
Standard macroeconomic models, from the classical IS-LM framework Hicks (1937) to modern New Keynesian Dynamic Stochastic General Equilibrium (DSGE) models Clarida et al. (1999); Woodford (2003), postulate that the economy is driven by three fundamental state variables:
  1. 1.
    Output Gap ($y_{t}$): Represents real economic activity and the business cycle (The IS curve relation).
  2. 2.
    Inflation ($\pi_{t}$): Represents the price level dynamics and aggregate supply constraints (The Phillips Curve).
  3. 3.
    Nominal Interest Rate ($i_{t}$): Represents the policy stance and liquidity preference (The Taylor Rule or LM relation).
This theoretical ”triad” suggests that the essential dynamics of the economy evolve on a low-dimensional manifold embedded in a 3-dimensional phase space $(y_{t},\pi_{t},i_{t})$. While a fourth dimension often identified as financial volatility or risk premia plays a critical role, it frequently acts as a regime-switching parameter that alters the curvature of the space rather than an independent state variable of the same order.
Therefore, extracting the first three principal components is not merely a data compression technique; it is an attempt to recover the axes of this latent macroeconomic manifold. We hypothesize that our 3-dimensional eigenportfolio embedding $\mathbf{X}_{t}$ spans the same subspace as these fundamental macro-variables, allowing us to interpret the geometric ”shape” of the market (e.g., the cyclical tori) as a direct manifestation of endogenous business cycles and policy feedback loops.
@@ 4.1 Ambient formulation and curvature drift
Brownian motion constrained to $M$ is the Stratonovich SDE (Hsu (2002a))
@@
where $B_{t}$ is standard Brownian motion in $\mathbb{R}^{3}$. In Ito form this becomes (Hsu (2002a))
@@ 4.2 Explicit SDEs by geometry
    The metric is diagonal: $g_{\theta\theta}=(R+r\cos\varphi)^{2}$, $g_{\varphi\varphi}=r^{2}$,do Carmo (1992), hence $g^{\theta\theta}=(R+r\cos\varphi)^{-2}$, $g^{\varphi\varphi}=r^{-2}$ and $\sqrt{|g|}=r(R+r\cos\varphi)$. The Ito SDE realizing generator $\tfrac{1}{2}\Delta_{g}$ is
@@
Then $g_{uu}=E(u)=a^{2}\sinh^{2}u+c^{2}\cosh^{2}u$, $g_{vv}=G(u)=a^{2}\cosh^{2}u$,do Carmo (1992), so $g^{uu}=1/E$, $g^{vv}=1/G$, $\sqrt{|g|}=\sqrt{E(u)G(u)}=a\cosh u\,\sqrt{E(u)}$. The Ito SDE (generator $\tfrac{1}{2}\Delta_{g}$) reads
@@ 4.3 Nonlinear forecasting via machine learning regressors
A Random Forest (Breiman, 2001) constructs an ensemble of $B$ regression trees $\{f_{b}(\cdot)\}_{b=1}^{B}$, each trained on a bootstrap resample of $\mathcal{D}$ and a random subset of predictors. The forecast corresponds to the ensemble mean:
@@ 4.4 Logarithmic and exponential mappings
Let $T_{x}M$ denote the tangent plane at $x\in M$.Pennec et al. (2019); do Carmo (1992)
@@
See Pennec et al. (2019); do Carmo (1992) for closed forms.
Torus $T^{2}(R,r)$ (angle chart).do Carmo (2016)
@@
Hyperbolic (hyperboloid chart).do Carmo (1992)
@@
Karcher mean and tangent PCA.Karcher (1977); Pennec et al. (2019)
@@ 4.5 Data-driven estimation of manifold parameters
A key ingredient of our pipeline is that the geometry is not treated as fixed; instead, its intrinsic parameters are inferred on the fly from the observed 3D path $\{x_{t}=(x_{t},y_{t},z_{t})\}_{t\leq t^{\star}}$. This subsection formalizes the estimation steps we implement for the sphere, torus and one–sheeted hyperboloid, matching the code used in our experiments.Cazals and Pouget (2005); Cohen-Steiner and Morvan (2006)
@@
For spherical segments, we first compute a Karcher (Fréchet) mean $\widehat{\mu}\in S^{2}$ of the points $\{x_{t}/\|x_{t}\|\}$ by iterating $\mu\leftarrow\exp_{\mu}\!\big(\tfrac{1}{n}\sum_{t}\log_{\mu}(x_{t})\big)$ until convergence;Karcher (1977)
@@
In practice, for stability we use $\widehat{R}$ from (25) while the instantaneous tube radius is proxied by the latest $\rho_{t^{\star}}$, which helps track slow deformations of the tube thickness in non–stationary segments. do Carmo (2016)
@@
On this chart, the log map again uses shortest arc wrapping on $v$ and simple differencing on $u$; the exp map re–adds those increments around the base point. do Carmo (1992)
@@ 4.6 Local Gaussian Curvature Based Geometry Inference
At each time $t$ we consider a window $\mathcal{W}_{t}=\{X_{s}:s\in[t_{0},t]\}$ ( expanding) with $|\mathcal{W}_{t}|\geq m_{0}$ points and compute a local quadratic Monge patch fit,Cazals and Pouget (2005); Cohen-Steiner and Morvan (2006),
@@
where $W_{t}=\mathrm{diag}(\,w_{1},\dots,w_{|\mathcal{W}_{t}|}\,)$ is optional exponential weighting (recent samples weighted more), with $w_{i}\propto e^{-\alpha(t-s_{i})}$. Given $\hat{a}_{t},\hat{b}_{t},\hat{c}_{t}$, the local Gaussian curvature at time $t$ is (Monge gauge approximation),Cazals and Pouget (2005),
@@
Towards validating that the method approximates satisfactorily the underlying Gaussian curvature, Fig. 2 recovers the expected signatures on benchmarks—-$K>0$ on $S^{2}$, mixed $K$ on $T^{2}$, and $K<0$ on $H^{2}$—and yields plausible, intermittent curvature on the finance path; “uniformly sampled” means draws from each surface’s Riemannian (area) measure (density $\propto dA$).do Carmo (1992)
@@
We form a Takens embedding, Takens (1981), over the window (Takens (1981)) $\mathcal{W}_{t}$:
@@
with delay $\tau$ and embedding dimension $m$. We compute Vietoris Rips persistent homology of $\mathcal{T}_{t}$ and count persistent 1-cycles (Betti $1$ features),Edelsbrunner and Harer (2010); Ghrist (2008). A torus satisfies $\beta_{1}=2$ (and $\beta_{2}=1$),Edelsbrunner and Harer (2010), so we flag
@@
with persistence threshold $\epsilon$ calibrated to the scale of the point cloud,Gidea and Katz (2018); Ismail et al. (2022); Arvanitis and Detsis (2024). In the current context, we implement the basic $\beta_{1}$–based detector over sliding windows.